Sammys Pizza opened on January 1, 2018. Sammys reported the following for cash revenues and cash expenses for the years 2018 to 2020 ash Revenues Cash Expenses 2819 512,098 2019 $22,8ee 2018 7,808 1e,eee $18,e0 2.8 2828 $32,888 Required a. What would Sammys Pizza report for net income and retained earnings for the years 2018, 2019, and 2020? c. Assume that Sammys Pizza paid a $3,000 dividend to stockholders in 2019 What would Sammys Pizza report for net income and retained earnings for 2019 and 2020
Year Net Income Earnings a. 2018 2019 2020 C. 2018 2019 2020
Answer #1

Answer:

Answer:

A )

Year

Net income

Retained
earning

1

5000

5000

2

12000

17000

3

14000

31000

C)

Year

Net income

Retained
earning

1

5000

5,000

2

12000

14,000

3

14000

28,000

Working notes for the above answer is as under

31 Cash Cash expanse Net Retained 32 Year revenue income earning C=A-B 12,0007, 34 35 36 37 7,000 22,000 10,000 2000 18,000 5,000 12,000 14,000 5,000 17,000 31,000 2 Cash Cash expanse Net Devidned Retained 38 Yea 39 income C=A-B revenue paid earnin E=C-D 12,0007, 7,000 22,000 10,000 2000 18,000 5,000 12,000 14,000 5,000 3,000 14,000 028,000 0 41 2 42

31 Cash Cash Retained 32 Year revenue expanses Net income earning A-B +134-J 34 =+135-J35 +136-J36 C 34 1 35 2 36 3 37 12000 7000 10000 18000 5000 +L34+K35 -+L35+K36 32000 Cash Cash Devidned 38 Year 39 40 1 41 2 42 3 Net income C=A-B =+140-140 =+41-41 -+142-J42 Retained earnin E-C-D 5000 revenue expanses pai 12000 7000 10000 18000 0 3000 0 +M40-K41-L41 32000 +M41+K42-L42